### v1

# Many-Task Federated Fine-Tuning via Unified Task Vectors

Abstract
Federated Learning (FL) traditionally assumes homogeneous client tasks; however, in real-world scenarios, clients often specialize in diverse tasks, introducing task heterogeneity. To address this challenge, Many-Task FL (MaT-FL) has emerged, enabling clients to collaborate effectively despite task diversity. Existing MaT-FL approaches rely on client grouping or personalized layers, requiring the server to manage individual models and failing to account for clients handling multiple tasks. We propose MaTU, a MaT-FL approach that enables joint learning of task vectors across clients, eliminating the need for clustering or client-specific weight storage at the server. Our method introduces a novel aggregation mechanism that determines task similarity based on the direction of clients’ task vectors and constructs a “unified” task vector encapsulating all tasks. To address task-specific requirements, we augment the “unified” task vector with lightweight modulators that facilitate knowledge transfer among related tasks while disentangling dissimilar ones. Evaluated across 30 datasets, MaTU achieves superior performance over state-of-the-art MaT-FL approaches, with results comparable to per-task fine-tuning, while delivering significant communication savings.
1 Introduction
Federated Learning (FL) has emerged as a collaborative learning paradigm, enabling joint training of neural network models across edge devices (refereed to as clients), while keeping data localized McMahan et al. (2017). Traditional FL settings assume clients work on the same set of tasks; yet, in real-world cross-silo FL the inherent task heterogeneity among clients can naturally occur as clients typically specialize in specific tasks. This phenomenon, known as task heterogeneity, represents a novel and under-explored form of heterogeneity in FL, adding complexity to the learning process.
Recent studies Chen et al. (2023); Zhuang et al. (2023); Cai et al. (2023) have underscored the significance of task heterogeneity in FL. To address this diversity, traditional FL has evolved into Many-Task Federated Learning (MaT-FL) Cai et al. (2023) to enable clients to effectively collaborate, despite specializing in different tasks. It is worth to mention that while “many-task” has been interchangeably used as “multi-task” in conventional ML, early Multi-Task Federated Learning (MTFL) mainly addressed personalized FL, a subset of MaT-FL. MaT-FL is thus used to distinguish it from prior MTFL works Cai et al. (2023); Muhamed et al. (2024); Zhuang et al. (2023). Current MaT-FL approaches address task heterogeneity by focusing on client grouping or split FL. For instance, FedBone Chen et al. (2023) employs split FL whereas the server needs to sustain a unique model per client. Conversely, MaT-FL Cai et al. (2023) and MAS Zhuang et al. (2023) concentrate on dynamically grouping client models based on task similarity, aggregating models among clients that handle similar tasks. Nevertheless, these strategies necessitate the server managing individual models for each client or task, while they do not consider cases in which each client hold more than one task. Nevertheless, these strategies require the server to manage individual models for each client or task and fail to address scenarios where clients hold more than one task (e.g., a self-driving car handling tasks like object detection, lane recognition, and pedestrian tracking simultaneously). Training a single model among tasks offers considerable benefits in federated settings. Training individual models for each task is resource-intensive, whereas a unified model not only conserves resources but also potentially enhances performance by leveraging learning from diverse auxiliary tasks. This capability is particularly vital in pragmatic FL scenarios in which data availability dramatically varies across tasks — some enjoying ample data, while others are data-starved.
Concurrently, the advent of large-scale deep models trained on massive amounts of data in a self-supervised fashion has ushered in a new era of Foundation Models (FMs). These models, exemplified by Large Language Models (LLMs) and vision Foundation Models (vFMs) hold the promise of adapting to a wide range of downstream tasks. Furthermore, Parameter Efficient Fine-Tuning (PEFT) techniques, allowing FMs to reach performance levels comparable with fully fine-tuned models with only minimal adjustments to their parameters, have recently seen immense popularity, with LoRA adapters standing out for their efficacy and broad applicability in FL Yi et al. (2023); Cho et al. (2023); Nguyen et al. (2024); Ping et al. (2024).
This paper addresses the challenge of adapting pretrained Foundation Models (FMs) to the MaT-FL setting, where clients hold an arbitrary (>$1$) number of tasks, and we aim to train a unified model across all tasks. We propose Many-Task FL via Unified Task Vectors (MaTU), a novel approach that eliminates the need for clustering or client-specific weight storage at the server, enabling natural knowledge transfer across tasks in FL. Inspired by recent advancements in model merging and Task Arithmetic (TA) Ilharco et al. (2023); Yadav et al. (2023a)MaTU introduces a task vector aggregation mechanism tailored for MaT-FL. Specifically, we determine task similarity directly based on the direction of clients’ task vectors and constructs a “unified” task vector that encapsulates all tasks. To handle task-specific requirements, we augment the “unified” vector with lightweight modulators (i.e., binary task-specific masks and scalers) that facilitate knowledge transfer between related tasks while promoting weight disentanglement for dissimilar ones, ensuring robust performance across heterogeneous tasks in FL. Concisely, our main contributions are as follows:
  • We introduce MaTU, a MaT-FL approach enabling joint training across clients which holding multiple tasks without requiring server-side clustering or client-specific weight storage.
  • We propose a novel task vector aggregation mechanism that constructs a unified task vector across tasks, while leveraging lightweight task-specific modulators to promote weight disentanglement and knowledge transfer.
  • We conduct a comprehensive performance evaluation across 30 datasets, where MaTU outperforms state-of-the-art MaT-FL regimes with performance comparable to per-task adaptation fine-tuning, all while delivering significant communication savings.
2 Related Work
Federated PEFT. Parameter Efficient Fine-Tuning (PEFT) has emerged as an alternative fine-tuning strategy, where most of the pre-trained model parameters are frozen, and only a small portion of task-specific parameters are trained. Among the various PEFT methods, the injection of additional adapter modules, fine-tuned independently from the pre-trained model — such as Low-Rank Adaptation (LoRA)Hu et al. (2021) — has achieved state-of-the-art results across many large pre-trained models.
In the context of FL, PEFT-based approaches have demonstrated success in enhancing privacySun et al. (2024), communication Tsouvalas et al. (2023), and computation Babakniya et al. (2023) efficiency. SLoRA Babakniya et al. (2023) addresses data heterogeneity in federated settings by utilizing multiple LoRAs, while FFA-LoRA Sun et al. (2024) enhances privacy in FL by applying differential privacy techniques to LoRA modules. Moreover, DeltaMask Tsouvalas et al. (2023) combines LoRAs with probabilistic encoding to achieve significant bitrate reductions during FMs fine-tuning in FL, and FS-LLM Kuang et al. (2023) extends PEFT to LLMs in the federated settings. Nonetheless, most PEFT-based FL approaches primarily focus on single-task settings, overlooking their potential in many-task settings. In our work, we explore combining LoRAs with model merging techniques, such as TIES Yadav et al. (2023b), to address task heterogeneity in FL, enabling the training of a unified model across multiple tasks while achieving significant reductions in communication and computation overhead.
MaT-FL. Many-Task Federated Learning (MaT-FL) has emerged to tackle task heterogeneity in FL. Although the term “many-task” mirrors “multi-task” in conventional ML, early Multi-Task Federated Learning (MTFL) primarily focused on personalized FL, a specific case of MaT-FL. Therefore, we adopt the term MaT-FL to distinguish it from earlier MTFL works, as seen in Cai et al. (2023); Muhamed et al. (2024); Zhuang et al. (2023). From MTFL approaches, FedProx Li et al. (2020), introduced a proximity term to limit client updates from deviating too far from the global model in MTFL. Alternatively, FedBone Chen et al. (2023) used split FL to introduce task-specific personalized layers to handle multiple tasks in FL.
Recent MaT-FL approaches primarily address task heterogeneity by focusing on client grouping Cai et al. (2023); Zhuang et al. (2023); Lu et al. (2023). MaT-FL Cai et al. (2023) and MAS Zhuang et al. (2023) dynamically group clients based on task similarity, enabling model aggregation among clients with similar tasks. FedHCA Lu et al. (2023) extends this by supporting a variable number of tasks across clients, creating a more flexible framework. More recently, NTKFedAvg Muhamed et al. (2024) introduced task arithmetic in MaT-FL, enabling clients to train task-specific adapters and leverage server-side adapter fusion for optimizing multiple tasks. To further improve task disentanglement during model aggregation, they applied Neural Tangent Kernel (NTK) linearization over the model prior to training. Nevertheless, group-based MaT-FL approaches require the server to manage multiple models for each task group, assume task relationships are known in advance, and introduce significant client-side overhead, making scalability a challenge as the number of tasks increases. Training a single model across tasks offers considerable benefits: it reduces resource consumption and can enhance performance by leveraging diverse auxiliary tasks. While NTKFedAvg attempts to address these challenges, it focuses on a single task per client, whereas in practice, clients may hold multiple tasks. MaTU bridges this gap by training a unified model across all clients and tasks, dynamically building task correlations to enable knowledge transfer, and using task vector aggregation based on task similarity, removing the need for multiple server-side models.
3 Methodology
Figure 1: MaTU’s training process: Clients update their task vectors, and transmit their “unified” task vector and task modulators to server, which constructs task masks, estimates task similarity via sign conflicts, and construct clients’ task vectors using the top-$\kappa$ similar tasks.
3.1 Preliminaries
Task Arithmetic. We denote with $\theta_{p}\in\mathbb{R}^{d}$ the pre-trained model weights of a FM. A task vector $\tau_{t}$ for task $t$ is defined as $\tau_{t}=\theta_{t}^{*}-\theta_{p}$, where $\theta_{t}^{*}$ refers to the fine-tuned model weights of task $t$. Essentially, $\tau_{t}$ can be seen as a representation of the direction (sign) and amount (magnitude) of movement relative to $\theta_{p}$ needed in the $d$-dimensional weights space to lead to a low loss region for the task $t$. Additionally, a scalar multiplication with $\lambda$ over $\tau_{t}$, i.e., $\tau_{new}^{t}=\lambda\cdot\tau_{t}$ yields models parameters along the training trajectory, with positive $\lambda$ values denoting task learning and negative $\lambda$ values indicating task negation (unlearning) Ortiz-Jimenez et al. (2023). Hence, adding $K$ task vectors, a many-task (“merged”) model can be obtained aiming to solve all $K$ tasks as follows:
$\theta_{\text{m}}=\theta_{p}+\sum_{k=1}^{K}\lambda_{k}\cdot\tau_{k}$ (1)
By only adjusting $\lambda_{k}$, the performance across $K$ tasks can be improved. Building on the observations from Ortiz-Jimenez et al. (2023), TIES Yadav et al. (2023a) demonstrated that resolving sign conflicts before model merging enhances weight disentanglement (i.e., separating the influence of different task vectors), leading to improved task performance. Furthermore, re-scaling task vectors often yields better results than using unmodified vectors Yang et al. (2024); Yu et al. (2024).
Unified Task Vector. Given a number of task vectors derived from a set of tasks, $t\in\mathcal{T}$, the “task unification” Huang et al. (2024) process begins by computing the aggregated sign vector across all task vectors, $\sigma$ = $\text{sgn}\left(\sum_{i=1}^{|\mathcal{T}|}\tau^{i}\right)$. Next, the “unified” task vector is as follows:
$\tau=\sigma\odot\mu~{},$ (2)
where $\mu$ refers to the magnitude vector constructed by extracting the maximum absolute value from the task vectors whose signs align with $\sigma$. This electing procedure has shown to reserve the maximum amplitude and sign information shared by the task vectors, thereby maximally reducing interference Huang et al. (2024).
Weight disentanglement. Weight disentanglement — the model’s ability to isolate the influence of different tasks on its weights — serves as a necessary condition for task arithmetic operations, and is an emerging property of pre-training Ortiz-Jimenez et al. (2023). Since the sign of each weight indicates the direction that minimizes loss for a given task Ortiz-Jimenez et al. (2023); Yadav et al. (2023a); Muhamed et al. (2024), resolving sign conflicts among task vectors can promote weight disentanglement, leading to improved individual task performance Yadav et al. (2023a). Therefore, in the context of the same pre-trained model, a high count of sign conflicts between two task vectors implies opposing influences on model weights, suggesting potential interference in their joint training.
3.2 MaTU :Efficient MaT-FL with Task Arithmetic.
Overview. Here, we present the general MaTU training pipeline (see Fig. 1). Clients initialize the same model $f$, parameterized by the pre-trained weights $\theta_{p}$, and train on their local datasets $D^{t}_{n}$ to generate individual task vectors $\tau^{t}_{n}$, which are combined into a unified task vector $\tau_{n}$. To account for the changes from $\tau^{t}_{n}$ to $\tau_{n}$, clients create lightweight modulators — binary masks $m^{i}_{n}$ and scaling values $\lambda^{i}_{n}$ — which are transmitted along with $\tau_{n}$. Upon receiving client updates, for each task, the server computes an average task mask, $\hat{m}^{t}$, to capture key areas in the task vectors, then performs task-specific aggregation to combine updates from clients sharing the task, followed by cross-task aggregation to enable knowledge transfer between related tasks. By averaging these two, the server creates the updated task vectors to compute unified task vectors and modulators for each client based on their assigned tasks, which then transmits to client. Next, clients use these modulators to adjust the unified task vector and derive the updated task vectors, marking the next federated round.
Notation. We use the following notation in the rest of the paper. Let $\mathcal{N}$ refer to a set of clients, each assigned a set of $k$ tasks $\mathcal{T}_{n}$ ($k_{n}=|\mathcal{T}_{n}|$), which may overlap across clients (i.e., the unique set of tasks is given by $\mathcal{T}=\bigcup_{n=1}^{|\mathcal{N}|}\mathcal{T}_{n}$). Each task $t\in\mathcal{T}_{n}$ is associated with a locally stored dataset $D^{t}_{n}$, containing $|D^{t}_{n}|$ samples. $A\in\{0,1\}^{N\times T}$ denotes a binary matrix representing task-to-client allocations, where $A(n,t)$=$1$ if client $n$ holds task $t$, and 00 otherwise. The neural network is denoted by $f$ and parameterized by weights $\theta$, where $\theta_{p}$ represents its pre-trained weights, and $\theta^{*}_{t}$ denotes its optimized weights after local training on task $t$. For each task $t$, its task vector is given by $\tau^{t}=\theta^{*}_{t}-\theta_{p}$, where $\sigma^{t}=\text{sgn}(\tau^{t})$ denotes the sign vector of $\tau_{t}$, indicating the direction of movement (positive or negative) relative to $\theta_{p}$.
Local Training with many-tasks. In a given round $r\in R$, the $n$-th client trains individually across the set of $k_{n}$ locally stored tasks, and derives a set of local task vectors, one for each task $t\in\mathcal{T}_{n}$. Next, the “task unification” process is performed to derive the client’s “unified” task vector, $\tau_{n}=\sigma_{n}\odot\mu_{n}$.
As a single task vector cannot capture all task-specific weights, leading to performance degradation, we introduce lightweight task-specific modulators that refine $\tau_{n}$ to approximate task-specific vectors, as in Huang et al. (2024). Specifically, we construct a set of task-specific masks, $\mathcal{M}_{n}=\{m^{i}_{n}\}^{k_{n}}_{i=1}$, where each mask is defined as $m^{i}_{n}=(\tau^{i}_{n}\odot\tau_{n}>0)$ for each task $t\in\mathcal{T}_{n}$; thus, aligning the direction of $\tau_{n}$ to each task vector. Similarly, to account for the magnitude shift between $\tau_{n}$ and individual task vectors, we introduce task-specific scaling parameters $\lambda_{n}=\left\{\lambda^{i}_{n}\right\}_{i=1}^{k_{n}}$, where $\lambda^{i}=\frac{\sum|\tau^{i}_{n}|}{\sum|m^{i}_{n}\odot\tau_{n}|}$.
Many-tasks Aggregation. Once the local training at round $r$ is completed, the server holds each client’s “unified” task vector and task-specific modulators (i.e., masks and re-scaling parameters). We aim to aggregate task vectors among clients holding a given task, while enabling dynamic knowledge transfer across similar tasks. Thus, we propose a simple-yet-effective many-tasks aggregation scheme tailored for MaT-FL, incorporating both task-specific and cross-task aggregation mechanisms.
Task-specific Aggregation. For each task $t$, we compute the average task-specific mask $\hat{m}^{t}$ across the clients that hold this task, defined as $\mathcal{N}^{t}=\{n\in\mathcal{N}\mid A(n,t)=1\}$, by performing the following element-wise operation on its elements:
$[\hat{m}^{t}]_{j}=\begin{cases}1&\text{if }\alpha^{t}_{j}\geq\rho\\ \alpha^{t}_{j}&\text{otherwise}\end{cases},~{}\text{where}~{}\alpha^{t}_{j}=\left|\frac{1}{|\mathcal{N}^{t}|}\sum\limits_{n\in\mathcal{N}^{t}}\text{sgn}(m^{t}_{n}\odot\tau_{n})\right|$
[…]
As demonstrated in Tenison et al. (2023), the agreement score $\alpha$ is closely related to client heterogeneity, with highly heterogeneity among clients resulting in $\alpha$ close to 00. As clients optimize their task vectors across distinct sets of tasks, the impact on each task-specific vector entry varies due to unique interference from their remaining tasks, exacerbating their heterogeneity. By adjusting the magnitude of updates based on the agreement score $\alpha$, we effectively reduce task interference, enhancing task separation; thus, promoting weight disentanglement during FL training stage. Using the $\hat{m}^{t}$, we then compute the average task-specific vectors across clients as:
$\hat{\tau}^{t}=\sum_{n\in\mathcal{N}^{t}}\gamma_{n}^{t}\cdot\lambda_{n}^{t}\cdot\hat{m}^{t}\odot\tau_{n}^{t}~{},$ (4)
where $\lambda_{n^{\prime}}^{t}$ is the task-specific re-scaler for client $n^{\prime}$ and task $t$, $\hat{m}_{t}$ is the aggregated task-specific mask for task $t$ (Eq. 3), and $\gamma_{n}^{t}$ represents clients’ weight in the aggregation based on their respective dataset size (i.e., $\gamma_{n}^{t}=|D^{t}_{n}|/\smash{\sum_{n\in\mathcal{N}^{t}}|D^{t}_{n}|}$), similar to FedAvg McMahan et al. (2017). In essence, $\hat{\tau}^{t}$ captures the task-specific information aggregated across clients in the FL process.
Cross-task Aggregation. Apart from capturing knowledge among clients holding task $t$, leveraging information from similar tasks can enhance task-specific performance, especially for data-scarce tasks Muhamed et al. (2024). To identify redundancies that facilitate joint learning and enhance individual task performance, we first examine task similarity—the relationships among tasks Zamir et al. (2018). Most task similarity metrics evaluate knowledge transfer benefits based on task-specific data Gholami et al. (2023); Bao et al. (2022) or labels Ding et al. (2022); yet, access to these resources is often limited in FL. Instead, as a high count of sign conflicts between task vectors for the same pre-trained model indicates potential interference in joint training Yadav et al. (2023a), we construct a task similarity matrix from sign conflicts among aggregated task-specific vectors. Specifically, the similarity between tasks $t$ and $t^{\prime}$ is defined as follows:
$S(t,t^{\prime})=\frac{1}{2}\left(\frac{1}{d}\sum\limits_{i=1}^{d}\left(\text{sgn}\left([\hat{\tau}^{t}]_{i}\right)\cdot\text{sgn}\left([\hat{\tau}^{t^{\prime}}]_{i}\right)\right)+1\right)$
[…]
$\tilde{\tau}^{t}=\sum_{t^{\prime}\in\mathcal{Z}^{t}}S\left(t,t^{\prime}\right)\cdot\hat{m}^{t}\odot\hat{\tau}^{t^{\prime}}~{},$ (6)
Note that we use $\hat{m}^{t}$, rather than $\hat{m}^{t^{\prime}}$, to adapt the $t$-th task’s vector based on the aggregated task vectors of task $t^{\prime}$, effectively enabling knowledge transfer across tasks.
Forming clients updates. Given Eq. 4 and 6, we can compute the aggregated task-specific vector for task $t$ on round $r$, as:
$\tau^{t,r+1}=\hat{\tau}^{t,r}+\tilde{\tau}^{t,r}=\underbrace{\sum_{n\in\mathcal{N}^{t},r}\gamma_{n}^{t}\cdot\lambda_{n}^{t,r}\cdot\hat{m}^{t,r}\odot\tau_{n}^{t,r}}_{\text{same-task}}~{}~{}+~{}~{}\underbrace{\sum_{t^{\prime}\in\mathcal{Z}^{t,r}}S\left(t,t^{\prime}\right)\cdot\hat{m}^{t,r}\odot\hat{\tau}^{t^{\prime},r}}_{\text{cross-tasks}}$
(a) Cosine Similarity
(b) LogME
(c) E-Trans
(d) WTE
(e) FIM
(f) Sign (Ours)
Figure 2: Comparison of sign vectors vs. state-of-the-art transferability estimation metrics across $8$ datasets. In all metrics except WTE, higher values correspond to higher correlation.
For the set of tasks assigned to each clients, $\mathcal{T}_{n}$, we can now utilize the newly computed aggregated task-specific vector to derive their updated “unified” task vector, $\tau^{r+1}_{n}$, and subsequently compute the set of task-specific masks, $\mathcal{M}^{r+1}_{n}=\{(\tau^{i,r+1}\odot\tau^{r+1}_{n}>0)\}^{k_{n}}_{i=1}$, and scaling parameters $\lambda^{r+1}_{n}=\left\{\frac{\sum|\tau^{i,r+1}_{n}|}{\sum|m^{i,r+1}_{n}\odot\tau^{r+1}_{n}|}\right\}_{i=1}^{k_{n}}$. The client’s “unified” task vector, $\tau^{r+1}_{n}$, and task-specific modules, $\mathcal{M}^{r+1}_{n}$ and $\lambda^{r+1}_{n}$, are then transmitted to the client-side for subsequent training. To train task $t$, the client must modulate the “unified” vector as $\dot{\tau}^{r+1}_{t}=\lambda^{t,r+1}_{n}\cdot m^{t,r+1}_{n}\odot\tau^{r+1}_{n}$, and injects it to the pre-trained model (i.e., $\theta^{r+1}_{t}=\theta_{p}+\dot{\tau}^{r+1}_{t}$). Note that, while it can be done on the server-side, performing the modulation locally on clients results in significant communication savings and is computationally cheap.
4 Performance evaluation
Datasets. We evaluate MaTU’s performance across $30$ vision classification datasets. Specifically, we consider $2$ model merging benchmarks: (i) an $8$-task benchmark comprising SUN397 Xiao et al. (2010), Cars Krause et al. (2013), RESISC45 Cheng et al. (2017), EuroSAT Helber et al. (2017), SVHN Netzer et al. (2011), GTSRB Stallkamp et al. (2011), MNIST LeCun et al. (2010), and DTD Cimpoi et al. (2014); and (ii) a large-scale $30$-task benchmark encompassing diverse vision tasks as proposed in Huang et al. (2024).
Baselines. We compare MaTU against centralized LoRA PEFT Hu et al. (2021), traditional FL methods (FedAvg McMahan et al. (2017), FedProx Li et al. (2020)), group-based approaches (MaT-FL Cai et al. (2023)), personalized FL approaches (FedPer Arivazhagan et al. (2019)) and related task-arithmetic methods like NTKFedAvg Muhamed et al. (2024). In FedPer, the last ViT block and classifier are used as personalized layers. We compare all baselines in terms of model performance (i.e., accuracy on test set) and client’s bitrate requirements (e.g. communicated bits per task in a round – $\mathrm{bpr}$). All experiments utilize ViT-B/32 Dosovitskiy et al. (2021) pre-trained on ImageNet-21k Deng et al. (2009) and employ LoRA Hu et al. (2021) with a rank of $16$ for PEFT, reporting test set accuracy.
FL Settings. We perform our FL simulations using Flower Beutel et al. (2020) with key parameters being number of clients ($N$), rounds ($R$), local epochs ($E$), clients’ participation rate ($\xi$), number of tasks ($T$), class concentration across clients ($\zeta_{c}$), and task concentration across clients ($\zeta_{t}$). We used a Dirichlet distribution, denoted as Dir($\alpha$), for both data and task splitting, where $\alpha$ represents the concentration parameter, following Li et al. (2021). We simulate non-IID settings with $\alpha$=$0.1$ ($\zeta_{c}$=$\zeta_{t}$=$0.1$) and train federated models for $100$ rounds ($R$=$100$) with $30$ clients ($N$=$30$), each performing one local training step per round ($E$=$1$). Note that for $\xi$<$1$, clients are randomly selected.
Figure 3: Pearson Correlation between sign-based and state-of-the-art task relatedness metrics across $8$ datasets.
Task Relatedness. We investigate the effectiveness of sign agreement among task vectors as a metric for measuring task similarity across datasets. To evaluate this, we conduct experiments with $8$ vision classification tasks, comparing our sign-conflict approach against well-established task similarity metrics, including cosine similarity of weights Vu et al. (2022), E-Trans Gholami et al. (2023), FIM Amari (1998), and WTE distance Liu et al. (2025). As shown in Fig.2, our sign-conflict method effectively identifies similar task groups among the datasets. Furthermore, Fig.3 highlights a strong Pearson correlation (>0.8) between the sign-conflict-based and all other task similarity metrics.
4.1 Results
Table 1: Performance evaluation of ViT-B/32 models on $8$ tasks in a single-task per client setting. Reported bitrates ($\mathrm{bpt}$) are in millions. Federated parameters are set to $N$=$30$, $E$=$1$, $\zeta_{t}$=$0.0$, $\xi$=$0.2$ and $R$=$100$ (except MaT-FL where $R$=$300$).
Methods SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg. Acc Bitrate ($\mathrm{bpt}$ M) $\downarrow$
Individual (Centralized) 74.81 76.93 96.05 99.43 97.05 98.57 99.67 79.21 90.21
FedAvg McMahan et al. (2017) 62.84 60.77 71.08 72.58 68.72 65.06 65.47 50.28 64.60 6.32
NTK-FedAvg Muhamed et al. (2024) 63.08 59.22 71.98 72.37 67.34 66.13 67.04 54.97 65.13 6.32
FedProx Li et al. (2020) 66.58 63.64 74.95 76.26 70.02 62.86 67.65 55.16 67.14 6.32
FedPer Arivazhagan et al. (2019) 68.57 69.39 88.37 90.33 90.28 91.53 92.44 67.98 82.11 5.36
MaT-FL Cai et al. (2023) 67.02 65.19 85.13 82.68 85.90 80.16 87.08 60.12 76.66 6.32
MaTU (Ours) 71.61 69.17 91.27 92.59 93.59 94.10 96.16 70.09 84.32 6.32
Table 2: Performance evaluation of ViT-B/32 models on $8$ tasks in a multiple-task per client setting. Reported bitrates ($\mathrm{bpt}$) are in millions. Federated parameters are set to $N$=$30$, $E$=$1$, $\zeta_{t}$=$0.5$, $\xi$=$0.2$ and $R$=$100$ (except MaT-FL where $R$=$300$).
Methods SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg. Acc Bitrate ($\mathrm{bpt}$ M) $\downarrow$
Individual (Centralized) 74.81 76.93 96.05 99.43 97.05 98.57 99.67 79.21 90.21
FedAvg McMahan et al. (2017) 50.21 48.50 56.97 57.86 55.19 52.04 52.26 40.01 51.63 20.43
NTK-FedAvg Muhamed et al. (2024) 55.37 53.27 62.30 63.62 58.56 52.79 56.27 46.28 56.06 20.43
FedProx Li et al. (2020) 57.64 53.84 63.26 64.44 59.17 52.90 56.76 42.54 56.32 20.43
FedPer Arivazhagan et al. (2019) 41.29 38.26 57.01 61.20 57.06 55.06 36.32 40.74 48.37 17.16
MaT-FL Cai et al. (2023) 61.52 56.67 76.48 73.95 76.96 69.67 74.64 50.88 67.60 20.43
MaTU (Ours) 66.40 63.46 86.36 87.26 88.15 88.51 91.08 64.69 79.47 8.04
Single-task Clients. We explore a “simplified” version of MaT-FL, where each client holds a single task. Specifically, we perform experiments with $30$ clients ($N$=$30$) under low client participation rate ($\xi$=$0.2$), ensuring no overlap among client’s tasks ($\zeta_{t}$=$0.0$) in the $8$-task benchmark. We report our findings in Table 1, from which we observe that MaTU achieves the highest average accuracy (84.32%) among all baselines – less than a 6% gap to individual task fine-tuning performance. Traditional FL methods like FedAvg, NTK-FedAvg, and FedProx, with average accuracies of 64.60%, 65.13%, and 67.14%, respectively, struggle to merge task-specific updates effectively, even in single-task settings. The Neural Tangent Kernel (NTK) linearization provides minimal improvements over standard FedAvg in task disentanglement (less than 1%). While MaT-FL improves performance by mitigating task conflicts through cosine similarity-based grouping aggregation, it exhibits the slowest convergence among all baselines due to restricted client interactions. FedPer’s task-specific personalization layers deliver strong performance (82.11%) in single-task settings – as they are explicitly tailored to individual tasks. In contrary, MaTU outperforms FedPer by 2.21% on average, requiring only a marginal increase in bitrate. These results highlight MaTU’s ability to mitigate task interference via task-specific masks in aggregation, ensuring task disentanglement and superior performance without added communication overhead compared to standard FL regimes.
Figure 4: Performance evaluation of ViT-B/32 models on the $30$-task benchmark. We report test set accuracy normalized to individual task fine-tuning performance. Federated parameters are set to $N$=$30$, $R$=$300$, $E$=$1$, $\zeta_{t}$=$0.2$, and $\xi$=$1.0$.
Multiple-task Clients. We now explore a generalized MaT-FL scenario, where each client can hold an arbitrary (>$1$) set of tasks. For this, we perform experiments with $30$ clients ($N$=$30$) in both $8$-task and $30$-task benchmarks. Here, task conflicts are expected to intensify both within and across clients, as no restrictions are placed on task assignments (i.e., task groups are formed randomly given $\zeta_{t}$), allowing conflicting tasks to exist on a client. We report our results for both benchmarks in Table 2 and Fig. 4.
$8$-task benchmark: Table 2 demonstrates that MaTU maintains strong performance in multiple-task settings, achieving an average accuracy of 79.47% – a modest 5% drop from single-task settings – while outperforming all baselines and effectively minimizing task interference via masking. Traditional FL methods, such as FedAvg and FedProx, exhibit significant performance deficits, with average accuracies well below their single-task counterparts, underscoring their limitations in addressing task heterogeneity. Notably, NTK-FedAvg surpasses these baselines, likely benefiting from improved task disentanglement through NTK-based linearization. In stark contrast, FedPer experiences a severe performance decline ($\approx$40%) when transitioning from single to multiple-task settings, reflecting its inability to manage intra-client task heterogeneity due to its reliance on personalization layers. This highlights that personalized FL approaches are not designed to address the broader challenges of task heterogeneity, underscoring that personalization represents a subset of the more general MaT-FL problem. Conversely, MaT-FL achieves higher performance by employing similarity-based grouping to mitigate task conflicts, reducing interactions among dissimilar clients and tasks, but still lags MaTU by 10%. We further analyze how MaT-FL and MaTU handle such conflicts in Fig.˜6.
Beyond superiority in terms of test set model accuracy, MaTU demonstrates exceptional communication efficiency in multiple-task settings, as illustrated by the $\mathrm{bpt}$ column in Table 2. Competing methods transmit multiple adapters — one for each task held by a client — resulting in significantly higher communication costs and limited scalability. In contrast,MaTU incurs minimal overhead by transmitting a single “unified” task vector per client, supplemented by compact task modulators (a binary mask and scalar value per task).
$30$-task benchmark: We evaluate MaTU’s scalability to $30$ tasks and compare it to MaT-FL, the strongest-performing baseline in multiple-task settings (see Table 2). As shown in Fig.4MaTU achieves an average normalized performance of 77.40%, significantly surpassing MaT-FL’s 52.62% across most datasets, except MangoLeafDB. This demonstrates MaTU’s ability to minimize task interference through training and delivers performance comparable to individual task fine-tuning. An elaborate analysis of MaTU’s effectiveness with respect of increase of client’s tasks is presented in Fig. 5. More importantly, beyond superior joint training performance, MaTU constructs a “unified” task vector that encapsulates all tasks and can be swiftly adapted to individual tasks using computationally lightweight modulators, enabling significant storage savings during inference, particularly as model size scale.
Scaling with more client tasks. We further analyze how MaTU handles an increasing number of tasks per client. Specifically, we conduct experiments on the $30$-task benchmark, varying each client’s task groups from $2$ to $30$, and report both (i) communication cost per round (in MB) and (ii) average normalized performance across tasks (vs individual task fine-tuning). As shown in Fig.5(a)MaTU introduces minimal communication overhead as client tasks scale. This efficiency arises from transmitting a single adapter per client, regardless of number of tasks, supplemented by lightweight task-specific modulators (a binary mask and a scalar value), unlike the multiple adapters required byMaT-FL and other baselines. Furthermore, in terms of model performance, Fig.5(b) shows that MaT-FL suffers from a sharp performance drop when clients are assigned more than $5$ tasks, highlighting its limitations in managing high task heterogeneity. In contrary, MaTU demonstrates to have a solid performance – even when clients train on over $15$ tasks – highlighting its ability to effectively disentangle tasks in the weight space through task-specific modulators.
(a) Comm. Overhead
(b) Norm. Performance.
Figure 5: Impact of scaling number of tasks assigned to clients on MaTU. Performance evaluation of ViT-B/32 on the $30$-task benchmark. We report (a) communication cost for 1 federated round, and (b) normalized test set accuracy vs individual task fine-tuning. Federated parameters are set to $N$=$10$, $R$=$300$, $\xi$=$1$ and $E$=$1$.
Highly Conflicting task groups. Here, we evaluate MaTU’s performance in scenarios with highly conflicting tasks within client’s task group. Using the $8$-task benchmark, we focus on the $3$ distinct task clusters identified in Fig.2. Specifically, we conduct experiments with $10$ clients, each assigned a fixed group of three tasks, designed to include no, $2$, or $3$ highly dissimilar tasks, referred to as no conflict, 2-conflict, and 3-conflict task groups, respectively. This setup enables us to assess how MaTU handles task conflicts on the clients and compare its performance against baselines. As shown in Fig. 6MaTU achieves consistently high performance with minimal drop (less than $6$%), unlike baselines such as FedPer, which struggles to train multi-task models even without conflicting task groups, and MaT-FL, which suffers from a more pronounced performance drop as task conflicts increase.
Figure 6: Impact of conflicting tasks on MaTU. Performance evaluation of ViT-B/32 on the $8$-task benchmark. We report normalized test set accuracy vs individual task fine-tuning. Federated parameters are set to $N$=$10$, $R$=$300$, $\xi$=$1$ and $E$=$1$.
5 Conclusions
In this work, we introduced MaTU, a MaT-FL approach designed to address task heterogeneity by relying on task vectors that creates a “unified” task vector – encapsulating all tasks – together with lightweight task modulators. By eliminating the need for clustering or client-specific weight storage, MaTU enables scalable and communication-efficient training in MaT-FL. Our evaluations across $30$ datasets demonstrated that MaTU not only achieves superior performance compared to state-of-the-art MaT-FL regimes but also delivers results comparable to per-task fine-tuning while significantly reducing communication overhead; thus highlighting MaTU’s potential under highly heterogeneous FL scenarios.